how muny pounds of dried fruiy costing 1.50$ per pound should be mixed with 12 pounds of granola that costs 2.00 per pound in order to make a mixture that costs 1.80$ per pound
Answer by Alan3354(69443)

Same problem, different numbers.
-----------------------------------
How many liters of a 65% acid solution must be mixed with 10% acid solution to get 320 liters of a 30% acid solution?
---------------
s = amount of 65%
t = amount of 10%
---
s + t = 320 (total solution)
65s + 10t = 320*30 (total acid)
==========
s + t = 320 --> t = 320 - s
65s + 10t = 320*30 = 9600
Sub for t
65s + 10(320-s) = 9600
75s + 3200 = 9600
s = 6400/75
s = 256/3 liters
t = 704/3 liters